# ThinkSound: Chain-of-Thought Reasoning in Multimodal Large Language Models for Audio Generation and Editing

### v2

@@ D.1 Details on Video-to-Audio Comparison
For the results in Table 1, we reproduce the results of Seeing and Hearing (Xing et al., 2024), V-AURA (Viertola et al., 2024), FoleyCrafter (Zhang et al., 2024), and MMAudio (Cheng et al., 2024a) using the official code and pre-trained models. For the other baselines, we use the generated samples provided by the authors, i.e., Frieren, V2A-Mapper, and Movie Gen (Polyak et al., 2024). Furthermore, the $\text{CLAP}_{\text{cap}}$ scores of MMAudio, MovieGen, and ThinkSound are 0.43, 0.44, and 0.49, respectively.
@@ Appendix E Limitation and Future
While the current MLLMs are capable of a strong understanding and reasoning of semantic information, they still have limitations in understanding the precise temporal and spatial information of video. For example, in the case of locating the exact timestamp of the sound event, MLLMs often fail to provide accurate results or provide wrong results. What’s more, the current open-source video-audio datasets for audio generation are limited in diversity and coverage, which may lack rare or culturally specific sound events. In the future, we will continue to explore more diverse and comprehensive datasets to improve the performance of our model. Furthermore, we will explore more effective methods to improve the temporal and spatial alignment of generated audio.

### v3

@@ D.1 Details on Video-to-Audio Comparison
For the results in Table 1, we reproduce the results of Seeing and Hearing (xing2024seeing), V-AURA (viertola2024temporally), FoleyCrafter (zhang2024foleycrafter), and MMAudio (cheng2024taming) using the official code and pre-trained models. For the other baselines, we use the generated samples provided by the authors, i.e., Frieren, V2A-Mapper, and Movie Gen (polyak2024movie). Furthermore, the $\text{CLAP}_{\text{cap}}$ scores of MMAudio, MovieGen, and ThinkSound are 0.43, 0.44, and 0.49, respectively.
@@ D.4 Performance Comparisons between coarse-grained and fine-grained CoT
D.5 Effectiveness of T5 Encoder for CoT Structure
A key assumption of our work is that the T5 encoder can effectively capture the logical structure within the CoT reasoning steps. We provide both a theoretical rationale and empirical validation for this.
Theoretical Rationale
While T5 is pre-trained on general web text (C4), its Transformer architecture is fundamentally designed to capture long-range dependencies and compositional structure. The reasoning steps in our CoT (e.g., "first do A, then B happens, which causes C") are expressed through natural language syntax and logical connectives. T5’s contextual embeddings are well-suited to capture these structural and causal cues. The goal is not for T5 to "understand" reasoning in a human sense, but to produce a rich, structured conditioning signal that the downstream diffusion transformer can effectively leverage.
Empirical Validation
To prove that the T5 encoder effectively utilizes the structure of the CoT, we conducted an ablation study comparing our full model against two degraded variants: (1) Tags Only, which provides the model with only extracted keywords (e.g., "car," "rain," "dog bark"), removing all reasoning structure; and (2) Randomized CoT, which randomly shuffles the sentences within the CoT, preserving all keywords but destroying the logical flow. As shown in Table 12, both ablations lead to a significant performance drop. This strongly indicates that the performance gain is attributable to the stepwise logical structure of the CoT—as encoded by T5—and not merely the presence of more keywords.
Table 12: Ablation study on the importance of CoT structure. We compare our full model against variants with only keywords (Tags Only) and shuffled reasoning sentences (Randomized CoT). Results show that the logical structure is critical for performance.
[…]
Randomized CoT 40.52 1.56 1.35 0.51 0.43
Tags Only 39.35 1.60 1.38 0.50 0.42
Ours (Ordered CoT) 34.56 1.52 1.32 0.46 0.46
D.6 Impact of CoT Verbosity on Generation Quality
To mitigate the risk of verbose or hallucinated CoT text, we employ a two-pronged strategy: (1) rigorous quality control during the creation of our AudioCoT training data (see Appendix A.2), and (2) direct constraints on CoT generation during inference (e.g., $\leq$ 3 sentences, $\leq$ 77 tokens). To empirically validate this approach, we conducted an ablation study quantifying the impact of verbosity. We removed our length constraints to generate "Overly Detailed" CoT prompts and compared them against our proposed "Concise CoT." The results in Table 13 demonstrate a clear performance degradation with verbose reasoning, confirming that our dual strategy is effective and that concise reasoning is critical for achieving high-quality results.
Table 13: Impact of CoT verbosity. We compare our standard concise CoT against an overly detailed version generated without length constraints. Verbose reasoning leads to a clear performance degradation.
[…]
Overly-detailed CoT 43.56 1.61 1.55 0.54 0.35
Concise CoT (Ours) 34.56 1.52 1.32 0.46 0.46
D.7 Dedicated MLLM Reasoning Evaluation
The quality of the CoT reasoning is the linchpin of our framework. To substantiate our claims, we conducted a comprehensive evaluation of the generated CoT quality using both expert human annotators and automated LLM-based metrics.
Human Evaluation
We randomly sampled 100 VGGSound test cases. Two expert annotators rated each generated CoT on a 0-1 scale across five dimensions: (1) multimodal integration, (2) specificity of audio details, (3) feasibility for audio generation, (4) logical consistency, and (5) brevity/format compliance. The final score is the sum (max 5).
LLM-Based Similarity Evaluation
For large-scale analysis, we compared the generated CoTs to ground-truth CoTs using GPT-4-nano. The model was prompted to score similarity on a 0–5 scale, focusing on reasoning structure, causal/temporal relationships, and object-sound associations.
Results
The results, presented in Table 14, provide strong empirical evidence that our fine-tuned MLLM (ThinkSound) greatly outperforms other powerful MLLMs in generating high-quality, structured CoT for video-to-audio reasoning.
Table 14: Evaluation of MLLM-generated CoT quality. Our model is compared against baselines using human evaluation and LLM-based similarity scoring, demonstrating superior performance in generating high-quality CoT.
Model Human Score (0-5) LLM CoT Similarity (0-5)
Qwen2.5-VL-7B 3.78 3.95
Qwen2-Audio-7B 3.82 4.09
ThinkSound (VideoLLaMA2) 4.13 4.31
@@ Appendix E Limitation and Future
While the current MLLMs are capable of a strong understanding and reasoning of semantic information, they still have limitations in understanding the precise temporal and spatial information of video. For example, in the case of locating the exact timestamp of the sound event, MLLMs often fail to provide accurate results or provide wrong results. Moreover, the current open-source video-audio datasets for audio generation are limited in diversity and coverage, which may lack rare or culturally specific sound events. In the future, we will continue to explore more diverse and comprehensive datasets to improve the performance of our model. Furthermore, we will explore more effective methods to improve the temporal and spatial alignment of generated audio.